OpenStudy (anonymous):

Eliminate the parameter. x = 5t, y = t + 8

OpenStudy (anonymous):

find t from one eq . and put in the second eq.

OpenStudy (anonymous):

In order to eliminate the parameter t, you are basically putting the y equation in terms of x. We can do this since we know that x = 5t. So what is t in terms of x? (Hint: divide both sides by the 5)

OpenStudy (anonymous):

so were rewriting the equation, but how ? @thechocoluver445

OpenStudy (anonymous):

Since you know that x = 5t, you also know that t = 5/x. You can substitute that 5/x in for t in the second equation, y = t + 8.

OpenStudy (anonymous):

(That's what my hint was referring to but I guess it was a little unclear, sorry)

OpenStudy (anonymous):

y=-8 ?

OpenStudy (anonymous):

t=y-8 put in other.

OpenStudy (anonymous):

y = 5x + 8 y = x divided by five + 8 y = 5x - 8 y = x divided by five - 8 these are my answer choicews

OpenStudy (anonymous):

@surjithayer, when they are asking you to get rid of the parameter they mean leave the equation in terms of y and x, and not t

OpenStudy (anonymous):

x=5(y-8)

OpenStudy (anonymous):



OpenStudy (anonymous):

The second option is correct, then. @Zaria602, do you understand now?

OpenStudy (anonymous):

Oops my drawing is wrong! It should read (x/5)!!

OpenStudy (anonymous):

t = x/5! I also messed that up in my replies above, so sorry!

OpenStudy (anonymous):

kinda i dont understand how you are switching the x and t

OpenStudy (anonymous):

So you're given the equation x =5t. If you divide both sides by 5, you get that \[x = 5t\] \[\frac{ x }{ 5 } = \frac{ 5t }{ 5 }\] \[\frac{ x }{ 5 } = t\]

OpenStudy (anonymous):

Then you have the equation y = t +8 Since \[\frac{ x }{ 5 } = t\] you can substitute that in for t. Thus, you get \[y = t +8\] \[y = \frac{ x }{ 5 } + 8\]

OpenStudy (anonymous):

Eliminate the parameter. x = t2 + 2, y = t2 - 4 so how would it work with this question

OpenStudy (anonymous):

The squares may seem a little scary, but you don't have to worry about taking the sqaure root or anything like that because t^2 shows up in the y equation too. Instead of finding t in terms of x, you find t^2 since it's in the y equation as well. So you do a similar substitution: \[x = t^2 + 2\] Subtract 2 from both sides. \[x -2 = t^2\] And then you substitute that x -2 in for t^2 in the equation y = t^2 -4. \[y = t^2 -4\] \[y = x-2 - 4\] You can simplify this to \[y = x - 6\]

OpenStudy (anonymous):

y = x - 6, x ≥ 1 y = x + 6, x ≥ 1 y = x2 - 6, x ≥ 1 y = x2 + 6, x ≥ 1 these are my answer choices

OpenStudy (anonymous):

So it would be the first choice. The x >_ 1 is because of the x being squared.